OpenStudy (barreraa):

Write an equation to model the given situation. The area of a rectangle is 70 square inches. The length of the rectangle is 3 inches longer than the width. (Points : 1) 3w = 70 w + 3 = 70 w + 3w = 70 w(w + 3) = 70

OpenStudy (hysusonic):

do you know the formula for the area of a rectangle?

OpenStudy (barreraa):

sort of

OpenStudy (hysusonic):

length times width in case you need a refresher

OpenStudy (barreraa):

k thanks

OpenStudy (hysusonic):

wait forget everything ive said up until this point sorry

OpenStudy (barreraa):

ok

OpenStudy (hysusonic):

what are the factors of 70?

OpenStudy (barreraa):

The factors of 70 are 1, 70, 2, 5, 7, 10, 14, 35, -1, -70, -2, -5, -7, -10, -14, and -35. 1*70=70 2*35=70 5*14=70 7*10=70 -1*-70=70 -2*-35=70 -5*-14=70 -7*-10=70

OpenStudy (hysusonic):

positive factors only please, it makes your life easier

OpenStudy (hysusonic):

now which of the factors have a difference of 3

OpenStudy (barreraa):

1, 70, 2, 5, 7, 10, 14, 35

OpenStudy (hysusonic):

now which of the factors have a difference of 3

OpenStudy (barreraa):

2, 5, 7, 10,

OpenStudy (hysusonic):

ok now which 2 of these factors muliply to 70

OpenStudy (texaschic101):

A = L * W A = 70 L = W + 3 70 = W(W + 3)